A resistance of 5ohms is connected in parallel with another of a 4ohm and the parallel arrangements is connected in series with a 2ohm resistor and an end of 20V. Calculate the potential difference and current across each resistor


1
Expert's answer
2021-03-09T15:27:58-0500

Let's first find the equivalent resistance of the parallel arrangements of resistors:


"\dfrac{1}{R_{eq,p}}=\dfrac{1}{R_1}+\dfrac{1}{R_2},""R_{eq,p}=\dfrac{R_1R_2}{R_1+R_2}=\dfrac{5\ \Omega\cdot4\ \Omega}{5\ \Omega+4\ \Omega}=2.22\ \Omega."

Let's find the equivalent resistance of this parallel arrangements of resistors and the resistor connected in series:


"R_{eq, s}=R_{eq, p}+R_3=2.22\ \Omega+2\ \Omega=4.22\ \Omega."

Then, we can find the current in the circuit from the Ohm's law:


"I=\dfrac{V}{R_{eq,s}}=\dfrac{20\ V}{4.22\ \Omega}=4.74\ A."

Since the current in the series circuit is the same, "I=I_3=I_{eq,p}=4.74\ A."

Then, we can find the potential difference on the resistor "R_3":


"V_3=I_3R_3=2\ \Omega\cdot4.74\ A=9.48\ V."

Then, we can find the potential difference on the resitor "R_{eq, p}":


"V_{eq,p}=IR_{eq,p}=4.74\ A\cdot2.22\ \Omega=10.52\ V."

In the parallel circuits the voltage is the same across all elements, therefore:


"V_{eq,p}=V_1=V_2=10.52\ V."

Then, we can find the currents across resistors "R_1" and "R_2" from the Ohm's law:


"I_1=\dfrac{V_1}{R_1}=\dfrac{10.52\ V}{5\ \Omega}=2.1\ A,""I_2=\dfrac{V_2}{R_2}=\dfrac{10.52\ V}{4\ \Omega}=2.63\ A."

Answer:

"V_1=V_2=10.52\ V, V_3=9.48\ V."

"I_1=2.1\ A, I_2=2.63\ A, I_3=4.74\ A."